OpenStudy (anonymous):

solve by completing the square: y^2+7y+12=0

OpenStudy (anonymous):

FIrst subtract 12 from both sides, this leaves us with y^2 + 7y = -12 Now add to both sides (7/2)^2 => 49/4, this leaves us with y^2 + 7y + 49/4= -12 + 49/4 so now we have =, (y + 7/2)^2 = 1/4 Take square root of both sides y + 7/2 = 1/2 multiply both sides by 2 2y + 7 = 1 subtract 7 from both sides 2y = -6 divide both sides by 2 y = -3